OpenStudy (anonymous):

What are the solutions to the following system of equations? y = x2 + 3x − 7 3x − y = −2

OpenStudy (anonymous):

are you familiar with substitution method?

OpenStudy (anonymous):

no not really

OpenStudy (anonymous):

well it is not that hard what you need to do is substitute the u with the second expression and solve for x so it would be \[3x -(x^2+3x-7) = 7 \] see what i did? i got the y expression and plugged it in to the second y

OpenStudy (anonymous):

yes what would you do next?

OpenStudy (anonymous):

simplify it by combining the like terms so the 3x cancels out \[-x^2+3x-3x+7-7= 0 \]so \[-x^2+0+0 = 0 \]\[-x^2 = 0 \]\[x^2 = 0 \] \[x = 0 \]to find y, plug the x value in to one of the equation and solve for y

OpenStudy (anonymous):

so would it be (3, 11) and (−3, −7) @helpme1.2

OpenStudy (anonymous):

hello @helpme1.2